Grade 12th passAlgebra

Write a Pythagorean triplet whose one member is.
16

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m = 16
⇒ m = 16/2
= 8
m2– 1 =82– 1
= 64–1
=63
m2+1=82+1
=64+1
=65
∴ (16, 63, 65) is aPythagorean triplet.

Thanks